# IFT-UAM/CSIC-25-85

### v1

@@ III.1 High-temperature and zero modes
The only way to have a metastable minimum at high temperature, via renormalizable interactions, is to exploit the term linear in $T$ in the effective potential for $\phi$, balanced against the term proportional to $T^{2}$.  [^3] Allowing for non-renormalizable interactions of dimension $d_{\Lambda}$, we can have other terms of size $T^{4}(T/\Lambda)^{d_{\Lambda}}$, where $\Lambda$ is the effective scale. For us, it is then very important to assess the robustness of the matching condition in eq. (12) for $a_{3}$. And this leads us to the worry introduced by the presence of bosonic zero modes. As well known, the zero modes can enter a strongly coupled regime even if their couplings in the four dimensional theory are perturbative, since the small parameter of the perturbative series at finite temperature is $\varepsilon\approx\pi\bar{\lambda}_{S}T/m$, using the $S_{i}$ self-quartic for the estimate. For arbitrarily small values of $m$, which in our case corresponds to background field value $\phi$ close to the origin, the perturbative expansion is clearly subject to large corrections, which likely need to be re-summed to all orders, if at all possible. In usual situations, the procedure of re-summation allows eventually for a well-behaved limit $m\to 0$ for the zero mode, for example by shifting it with the 1-loop thermal self-energy Arnold:1992rz ; Parwani:1991gq (see also Curtin:2016urg ; Bittar:2025lcr for more recent updates).
@@
In practice, we want to simply avoid the need for re-summation, that is, we always work in the limit $\varepsilon\ll 1$. At our benefit, let us recall that $\phi$ is just a background field, so for us it is enough to explore field value of $\phi$ such that $\varepsilon$ is small, while still expanding the $J$-function for $T\gg m$. Then, it will be important to consistently check that we are in a situation of thick-wall phase transition where we do not need to explore the region close to the origin, or not even exactly know the potential there, but rather be in a limit where $\phi\gg T$. See also Buen-Abad:2023hex for similar considerations. Fortunately, this condition is achieved with a large $a_{2}$, which has at least the virtue of growing with $N$.
@@ IV Superheating and critical temperature
Indeed there will be a range of temperatures, during the heating, where we cannot at all neglect the instability of eq. (1). Therefore, we are led to see what are the implications of the mass instability, as this is strongly connected with the existence of a critical temperature, where two degenerate minima appears. Such a mass term renders the all scenario realistic and feasible.
Practically, the presence of the instability, can be traced back into all our formulas simply shifting $a_{2}\to a_{2}(T)=a_{2}-M^{2}/T^{2}$. And notice that the conditions of eq.s (3) and (4) can still apply with $a_{2}(T)$ in place of $a_{2}$.
However, the instability is a major deformation of the potential and of the dynamics, as it introduces a temperature $T_{0}$ defined by $a_{2}(T_{0})=0$, which identifies the alleged second-order phase transition end point, $T_{0}=M/\sqrt{a_{2}}$.
At the same time, now it is also possible to identify a critical temperature $T_{c}$, where the minimum in the origin and the one in $\phi_{*}(T_{c})$ are degenerate. The condition is $a_{2}(T_{c})=\frac{2}{9}a_{3}^{2}/a_{4}$, which reads off as
@@
It is interesting now to follow the time (or temperature) dependence of the action, since now there is a scale in the system. For all temperatures around $T\gtrsim T_{c}$, we expect a modification from the second line of eq. (6), induced by $a_{2}(T)$. However, for $T\gg T_{c}$ we should recover the scaling limit. Notice, importantly, that the conditions for the existence of the second minimum above the origin are now $T$-dependent and can be modified. In particular we can trade $M$ for $T_{c}$ with the above relation. For $T$ just marginally bigger than $T_{c}$, we have that $S_{3}/T$ is very large (exactly infinite at the critical temperature), eventually at high-$T$ it will reach the expression of (6). The question here is if there is a minimum for the action at finite $T$, or if the minimum is attained asymptotically at very high-temperatures. In the latter case, the most favourable condition for tunnelling is at high temperature.
@@ IV.1 Heating in the meta-stable minimum
By construction, we see that $T_{0}$ is always smaller than $M_{\phi}$, which suggest that thermal corrections due to a thermal population of $\phi$ are Boltzmann suppressed. Indeed we did not include them at all in the effective potential for $\phi$ in eqs. (2) and (12). As discussed, are the spectators $S_{i}$ that thermalise in the background of $\phi$. For this reason, it is just important to check the condition
@@ IV.2 Spinodal temperature
So far, we have outlined super-heating in the case where the existence of the second minimum at high temperature is set by conditions that hold for arbitrarily high temperatures. This leads to maximal super-heating up to the available $T_{\rm max}$, whose magnitude depends on the reheating process (see next section). We notice however, that for some choice of parameters super-heating can last indefinitely. This means that the system - after the temperature eventually drops below $T_{\rm max}$ - will cool down never restoring the symmetry at $\phi=0$.
To avoid this, we can relax the conditions for the existence of the barrier and minima of (3) and (4). In particular while holding on (4) with $a_{2}(T)$, we can just requires $\Delta$ just to be positive up to some temperature $T_{\rm SP}$ where it becomes zero. At this point, the minimum $\phi_{*}(T_{\rm SP})$ can be a saddle point or it can just disappear. This brings to the condition $a_{2}(T_{\rm SP})=\frac{1}{4}a_{3}^{2}/a_{4}$, which reads off as
@@
We notice that it is defined as soon as $\Delta<0$, i.e. $a_{3}^{2}<4a_{2}a_{4}$.
@@ IV.3 Models for the reheating sector
Let us finally comment on how the temperature can rise, since the decrease is easy to obtain thanks to the expansion of the Universe. We imagine that the sector described by $\phi$ and $S_{i}$ is populated by extracting energy from a “reheating sector”, $R$, with large energy density that is passed onto our sector with a rate $\Gamma_{R}$. At the same time, in the reheated sector, there is rate of thermalization, $\Gamma_{T}$, which has to be the fastest scale. Such a rate converts into thermal energy, at a temperature $T$, all the bits of energy extracted from the reservoir $R$. Our discussion here adheres to the one in Buen-Abad:2023hex .
@@
It is important, to even talk about the super-heating phase, that the system of $S_{i}$ thermalises below $T_{0}$, and not that it is immediately burnt at $T_{\rm max}$, otherwise we cannot argue in favour of a smooth growth from $f$ to $\phi_{*}(T)$. For this reason we want to focus on a scenario where $\Gamma_{R}\ll H$, such that the energy density of $R$ is transferred up to a time interval equal to $\Gamma_{R}^{-1}$. This way, the energy density of $S_{i}$ is growing in time up to when all $R$ has decayed. This will have an impact on the gravitational wave signal (see section VI).
@@
If all these conditions are met, on time scales smaller than $1/\Gamma_{R}$, the energy density of $S_{i}$ grows linearly with time $t\ll 1/\Gamma_{R}$, and the temperature raises as $T(t)\propto(t\Gamma_{R})^{1/4}\sqrt{H_{*}M_{\mathrm{Pl}}}/N^{\frac{1}{4}}$, where $M_{\mathrm{Pl}}$ is the Planck scale and $H_{*}$ is the Hubble scale at the onset of the heating phase. There is a short window where the temperature is raising, which we exploit next. We can think of this as the reheating phase of our Universe mentioned in the Introduction. This way $H_{*}$ would be the typical value of the Hubble scale after inflation during a matter dominated era, while the inflaton or any other reheating field is dominating the energy density prior to decay into our sector (and the SM).
@@ IV.4 Low-temperature phase transition below $T_{c}$
For all cases where nucleation happened before the system cools below $T_{c}$, the system will find itself in the origin while the temperature is decreasing adiabatically. At this stage, we can then ask whether the system will tunnel to the minimum away from the origin when the temperature drops back below $T_{c}$. Notice, that here we face an even more worrying aspect of the dynamics of the zero modes already discussed in section III.1. Indeed, close to $T_{0}$, the effective mass of the zero mode is extremely small compared to the temperature, making unreliable all computations that need to know precisely the effective potential at the origin.
Therefore, in section IV.5, we refrain to derive strong conclusions in this regime, except that in the case where there is a clear separation $T_{c}\gg T_{0}$ and the tunnelling proceeds in the thick-wall limit, where the details of the potential are irrelevant. In all cases where nucleation happens well before $T_{0}$ we can argue in favour of another first order phase transition.
@@ IV.5 Nucleation temperature
Figure 4: Illustration of the model with two phase transitions: while heating (top row) and cooling (bottom row). Heating is represented in red, while cooling is in blue. Left column: Schematic trajectories of the order‐parameter value. Solid curves show the actual evolution, arrows on the lines indicate the direction of time, and stars mark the nucleation temperatures for the inverse ($\leftarrow$) and direct ($\rightarrow$) transitions. Dashed extensions depict the hypothetical path if tunnelling did not occur. The red shaded area at the bottom indicates the thermalisation of the system below $T_{0}$. Right column: Corresponding bounce action $S_{3}/T$ (black curve) with the nucleation criterion from Eq. (18) overlaid (coloured lines for different Hubble rates). Solid segments represent the realised evolution. In both heating and cooling cycles, once the inverse transition occurs, the direct transition inevitably follows.
@@
For all practical conditions, the Hubble scale is constant while the system is heating up. Therefore the nucleation condition, during heating, is given by
@@
If a $T_{n}^{\leftarrow,h}$ exists, the system can even tunnel back to $\langle\phi\rangle\approx f$ when, after $T_{\rm max}$, the temperature cools adiabatically with the expansion of the universe. In that case, assuming radiation dominance, the Hubble scale is controlled by the SM temperature, which can differ by a factor of $1/\xi$ compared to $T$, usually $\xi\ll 1$. In that case, $S_{3}/T$ has to be compared with $\approx 4\log M_{\mathrm{Pl}}/T$, and there might be a case where $T_{n}^{\rightarrow,c}$ exists well before $T_{0}$. We require $T_{n}^{\rightarrow,c}\gg T_{0}$ to make reliable estimates based on the shape of the potential just away from the origin. It is amusing to see that there are or zero or two phase transitions, but not just one. Often, the two are also both strongly first order. We show in figure 4-left a region of parameter space where both $T_{n}^{\leftarrow,h}$ and $T^{\rightarrow,c}$ exist.
Else, if a $T_{n}^{\leftarrow,h}$ does not exist, most likely we have ended up in a region of maximal superheating where $S_{3}/T$ approaches a constant (see figure 2). In this case the action is just too big, since $4\log T/H_{R}$ just falls short to activate nucleation before $T_{\rm max}$. We notice, and this was not present in Buen-Abad:2023hex , that when the temperature drops below $T_{\rm max}$, the condition for nucleation compares $S_{3}/T$ almost constant and $4\ln(T/H)$, which grows as temperature drops. This implies the possible existence of a solution for a new nucleation can happen with a $T_{n}^{\leftarrow,c}$ while cooling. Eventually there will be another transition $T_{n}^{\rightarrow,c}$ to the minimum in $f$ while cooling. Again, or zero or two phase transitions. We show this parameter space in figure 4-right.
As a last comment, we notice that in the region of parameter space where a finite spinodal temperature $T_{\rm SP}$ exists, the inverse transition during heating is guaranteed to occur, and the corresponding direct transition on cooling inevitably follows. However, because the direct cooling transition nucleates at temperatures close to $T_{0}$, the potential barrier lies too much near the origin where perturbation theory is not under control; consequently, at sufficiently low temperatures the exact shape of the potential near $\phi=0$ is uncertain and one cannot definitively determine the order of this second transition.
@@ V Superheating: inverse transition
The phase transition at high temperature brings the system to a vacuum with a largeN-N number of thermal states. Since they are massless in the new vacuum, they gain momentum when the bubbles expand. Therefore the leading order contribution from the pressure onto the expanding bubble wall has opposite sign to standard transitions where the tunnelling is towards a minimum with lesser relativistic states Buen-Abad:2023hex . It has been realized recently Barni:2024lkj ; Barni:2025mud that this aspect is connected with more generic properties of the hydrodynamics of the plasma surrounding the expanding bubbles, which is treated as a perfect fluid affecting the dynamics of the transition and production of gravitational waves Espinosa:2010hh ; Caprini:2019egz ; Hindmarsh:2015qta ; Hindmarsh:2017gnf ; Giese:2020rtr ; Giese:2020znk ; Wang:2021dwl ; Ajmi:2022nmq ; Tenkanen:2022tly ; Wang:2022lyd ; Wang:2023jto .
@@
Here, instead the dynamics is as in Barni:2024lkj ; Barni:2025mud , where the bubbles still nucleate and expand much as in the standard case, but the plasma is flowing inward. In these transitions, the latent heat is negative, and the whole process is endothermic. They are called inverse.
@@
The hallmark of an inverse first‐order transition is that the latent heat is negative, which in the frame of the centre of the bubble corresponds to an inward fluid velocity into the newly nucleated region. A simple diagnostic of this behaviour is the generalised pseudo‐trace of the energy–momentum tensor Barni:2025mud ,
$\alpha_{\vartheta}\equiv{D\vartheta\over 3w_{+}}=\frac{4}{3\,w_{+}}\left[De-\frac{\delta e}{\delta p}Dp\right],$ (19)
@@
and this condition remains valid regardless of the specific equation of state or level of approximation. The quantity $\alpha_{\vartheta}$ is nothing but the thermodynamic latent heat, $L=D\vartheta$, computed away from the critical temperature, normalised with the enthalpy of the old phase.
We avoid a full hydrodynamic treatment in favour of the pseudo‐trace criterion, valid when temperature jumps across the wall are small. Recalling that $\bar{\theta}=e-p/c_{s,-}^{2}$ and $\alpha_{\bar{\theta}}=D\bar{\theta}/4w_{+}$ (see Giese:2020rtr ), one finds that the sign of $\alpha_{\bar{\theta}}$ uniquely identifies the transition type.
Figure 5 displays the allowed region in nucleation temperature. When $T_{n}^{\leftarrow}>T_{c}$, these are heating PTs, the latent heat is negative ($\alpha_{\bar{\theta}}^{\leftarrow}<0$), corresponding to inverse transitions—whether the barrier eventually vanishes at a finite spinodal ($T_{c}/T_{0}<3$) or persists indefinitely ($T_{c}/T_{0}>3$). Conversely, for $T_{n}^{\rightarrow}<T_{c}$ the cooling transitions feature $\alpha_{\bar{\theta}}^{\rightarrow}>0$ and exhibit the usual direct behaviour. In this last estimates we considered $a_{0,2,3}\gg a_{4}$ so that $c_{s}^{2}\simeq 1/3$, appropriate for theories with many relativistic degrees of freedom (cf. Sec. III). The solid red (blue) line indicating the spinodal ($T=T_{0}$) temperature can be regarded as the maximal superheating (supercooling) we can achieve.

### v2

@@ III.1 High-temperature and zero modes
We treat $\phi$ as a background field and not as a fully thermalized degree of freedom. This choice allows us to isolate in a controlled way the thermal origin of the barrier responsible for the super-heating mechanism. In our framework, the barrier arises from the non-analytic high-temperature contribution of the thermalized spectator bosons, which generates an effective cubic term $\sim\phi^{3}T$ in the finite-temperature potential.
For this reason, we consider $\phi$ as weakly coupled to a thermalized sector, acting as an order parameter probing the thermal structure generated by the bath fields. This assumption is not merely a simplifying choice, but is essential to cleanly identify and characterise the super-heating mechanism studied in this work.
In this regard, the only way to have a metastable minimum at high temperature, via renormalizable interactions, is to exploit the term linear in $T$ in the effective potential for $\phi$, balanced against the term proportional to $T^{2}$.  [^5] Allowing for non-renormalizable interactions of dimension $d_{\Lambda}$, we can have other terms of size $T^{4}(T/\Lambda)^{d_{\Lambda}}$, where $\Lambda$ is the effective scale. For us, it is then very important to assess the robustness of the matching condition in eq. (12) for $a_{3}$. And this leads us to the worry introduced by the presence of bosonic zero modes. As is well known, the zero modes can enter a strongly coupled regime even if their couplings in the four dimensional theory are perturbative, since the small parameter of the perturbative series at finite temperature is $\varepsilon\approx\pi\bar{\lambda}_{S}T/m$, using the $S_{i}$ self-quartic for the estimate. For arbitrarily small values of $m$, which in our case corresponds to background field value $\phi$ close to the origin, the perturbative expansion is clearly subject to large corrections, which likely need to be re-summed to all orders, if at all possible. In usual situations, the procedure of re-summation allows eventually for a well-behaved limit $m\to 0$ for the zero mode, for example by shifting it with the 1-loop thermal self-energy Arnold and Espinosa (1993); Parwani (1992) (see also Curtin et al. (2018); Bittar et al. (2025) for more recent updates).
@@
In practice, we want to simply avoid the need for re-summation, that is, we always work in the limit $\varepsilon\ll 1$. At our benefit, let us recall that $\phi$ is just a background field, so for us it is enough to explore field value of $\phi$ such that $\varepsilon$ is small, while still expanding the $J$-function for $T\gg m$. Then, it will be important to consistently check that we are in a situation of thick-wall phase transition where we do not need to explore the region close to the origin, or not even exactly know the potential there, but rather be in a limit where $\phi\gg T$. See also Buen-Abad et al. (2023) for similar considerations. Fortunately, this condition is achieved with a large $a_{2}$, which has at least the virtue of growing with $N$.
@@ IV Superheating and critical temperature
Indeed, there will be a range of temperatures, during the heating, where we cannot at all neglect the instability of eq. (1). Therefore, we are led to see what are the implications of the mass instability, as this is strongly connected with the existence of a critical temperature, where two degenerate minima appear. Such a mass term renders all scenarios realistic and feasible.
Practically, the presence of the instability, can be traced back to all our formulas simply shifting $a_{2}\to a_{2}(T)=a_{2}-M^{2}/T^{2}$. And notice that the conditions of eq.s (3) and (4) can still apply with $a_{2}(T)$ in place of $a_{2}$.
However, the instability is a major deformation of the potential and of the dynamics, as it introduces a temperature $T_{0}$ defined by $a_{2}(T_{0})=0$, which identifies the alleged second-order phase transition end point, $T_{0}=M/\sqrt{a_{2}}$, see blue curve in fig. 1.
At the same time, now it is also possible to identify a critical temperature $T_{c}$, where the minimum in the origin and the one in $\phi_{*}(T_{c})$ are degenerate, see black dashed curve in fig. 1. The condition is $a_{2}(T_{c})=\frac{2}{9}a_{3}^{2}/a_{4}$, which reads off as
@@
It is interesting now to follow the time (or temperature) dependence of the action, since now there is a scale in the system. For all temperatures around $T\gtrsim T_{c}$, we expect a modification from the second line of eq. (6), induced by $a_{2}(T)$. However, for $T\gg T_{c}$ we should recover the scaling limit. Notice, importantly, that the conditions for the existence of the second minimum above the origin are now $T$-dependent and can be modified. In particular we can trade $M$ for $T_{c}$ with the above relation. For $T$ just marginally bigger than $T_{c}$, we have that $S_{3}/T$ is very large (exactly infinite at the critical temperature), eventually at high-$T$ it will reach the expression of (6). The question here is if there is a minimum for the action at finite $T$, or if the minimum is attained asymptotically at very high-temperatures. In the latter case, the most favourable condition for nucleation is at high temperature.
@@ IV.1 Heating in the meta-stable minimum
By construction, we see that $T_{0}$ is always smaller than $M_{\phi}$, which suggest that thermal corrections due to a thermal population of $\phi$ are Boltzmann suppressed. Indeed we did not include them at all in the effective potential for $\phi$ in eqs. (2) and (12). As discussed, it is the spectator fields $S_{i}$ that thermalise in the background of $\phi$. For this reason, it is just important to check the condition
@@ IV.2 Spinodal temperature
So far, we have outlined super-heating in the case where the existence of the second minimum at high temperature is set by conditions that hold for arbitrarily high temperatures. This leads to maximal super-heating up to the available $T_{\rm max}$, whose magnitude depends on the reheating process (see next section). We notice, however, that for some choice of parameters, super-heating can last indefinitely. This means that the system - after the temperature eventually drops below $T_{\rm max}$ - will cool down never restoring the symmetry at $\phi=0$.
To avoid this, we can relax the conditions for the existence of the barrier and minima of (3) and (4). In particular while holding on (4) with $a_{2}(T)$, we can just require $\Delta$ just to be positive up to some temperature $T_{\rm SP}$ where it becomes zero. At this point, the minimum $\phi_{*}(T_{\rm SP})$ can be a saddle point or it can just disappear. This brings to the condition $a_{2}(T_{\rm SP})=\frac{1}{4}a_{3}^{2}/a_{4}$, which reads off as
@@
We notice that it is defined as soon as $\Delta<0$, i.e. $a_{3}^{2}<4a_{2}a_{4}$. Therefore, at the spinodal temperature, the barrier is disappearing, and the minimum away from the origin ceases to exist, as one can see from the red solid curve in fig. 1.
@@ IV.3 Models for the reheating sector
Let us finally comment on how the temperature can rise, since the decrease is easy to obtain thanks to the expansion of the Universe. We imagine that the sector described by $\phi$ and $S_{i}$ is populated by extracting energy from a “reheating sector”, $R$, with large energy density that is passed onto our sector with a rate $\Gamma_{R}$. At the same time, in the reheated sector, there is rate of thermalization, $\Gamma_{T}$, which has to be the fastest scale. Such a rate converts into thermal energy, at a temperature $T$, all the bits of energy extracted from the reservoir $R$. Our discussion here adheres to the one in Buen-Abad et al. (2023).
@@
It is important to even talk about the super-heating phase, that the system of $S_{i}$ thermalises below $T_{0}$, and not that it is immediately burnt at $T_{\rm max}$, otherwise we cannot argue in favour of a smooth growth from $f$ to $\phi_{*}(T)$. For this reason we want to focus on a scenario where $\Gamma_{R}\ll H$, such that the energy density of $R$ is transferred up to a time interval equal to $\Gamma_{R}^{-1}$. This way, the energy density of $S_{i}$ is growing in time up to when all $R$ has decayed. This will have an impact on the gravitational wave signal (see section VI).
@@
If all these conditions are met, on time scales smaller than $1/\Gamma_{R}$, the energy density of $S_{i}$ grows linearly with time $t\ll 1/\Gamma_{R}$, and the temperature raises as $T(t)\propto(t\Gamma_{R})^{1/4}\sqrt{H_{*}M_{\mathrm{Pl}}}/N^{\frac{1}{4}}$, where $M_{\mathrm{Pl}}$ is the Planck scale and $H_{*}$ is the Hubble scale at the onset of the heating phase. There is a short window where the temperature is rising, which we exploit next. We can think of this as the reheating phase of our Universe mentioned in the Introduction. This way $H_{*}$ would be the typical value of the Hubble scale after inflation during a matter dominated era, while the inflaton or any other reheating field is dominating the energy density prior to decay into our sector (and the SM).
@@ IV.4 Low-temperature phase transition below $T_{c}$
For all cases where nucleation happened before the system cools below $T_{c}$, the system will find itself in the origin while the temperature is decreasing adiabatically. At this stage, we can then ask whether the system will transition to the minimum away from the origin when the temperature drops back below $T_{c}$. Notice, that here we face an even more worrying aspect of the dynamics of the zero modes already discussed in section III.1. Indeed, close to $T_{0}$, the effective mass of the zero mode is extremely small compared to the temperature, making unreliable all computations that need to know precisely the effective potential at the origin.
Therefore, in section IV.5, we refrain from deriving strong conclusions in this regime, except that in the case where there is a clear separation $T_{c}\gg T_{0}$ and the nucleation proceeds in the thick-wall limit, where the details of the potential are irrelevant. In all cases where nucleation happens well before $T_{0}$ we can argue in favour of another first order phase transition.
@@ IV.5 Nucleation temperature
Figure 5: Illustration of the model with two phase transitions: while heating (top row) and cooling (bottom row). Heating is represented in red, while cooling is in blue. Left column: Schematic trajectories of the order‐parameter value. Solid curves show the actual evolution, arrows on the lines indicate the direction of time, and stars mark the nucleation temperatures for the inverse ($\leftarrow$) and direct ($\rightarrow$) transitions. Dashed extensions depict the hypothetical path if nucleation did not occur. The red shaded area at the bottom indicates the thermalisation of the system below $T_{0}$. Right column: Corresponding bounce action $S_{3}/T$ (black curve) with the nucleation criterion from Eq. (18) overlaid (coloured lines for different Hubble rates). Solid segments represent the realised evolution. In both heating and cooling cycles, once the inverse transition occurs, the direct transition inevitably follows.
@@
For all practical conditions, the Hubble scale is constant while the system is heating up. Therefore, the nucleation condition, during heating, is given by
@@
If a $T_{n}^{\leftarrow,h}$ exists, the system can even transition back to $\langle\phi\rangle\approx f$ when, after $T_{\rm max}$, the temperature cools adiabatically with the expansion of the universe. In that case, assuming radiation dominance, the Hubble scale is controlled by the SM temperature, which can differ by a factor of $1/\xi$ compared to $T$, usually $\xi\ll 1$. In that case, $S_{3}/T$ has to be compared with $\approx 4\log M_{\mathrm{Pl}}/T$, and there might be a case where $T_{n}^{\rightarrow,c}$ exists well before $T_{0}$. We require $T_{n}^{\rightarrow,c}\gg T_{0}$ to make reliable estimates based on the shape of the potential just away from the origin. It is amusing to see that there are either zero or two phase transitions, but not just one. Often, the two are also both strongly first order. We show in figure 5-left a region of parameter space where both $T_{n}^{\leftarrow,h}$ and $T^{\rightarrow,c}$ exist.
Else, if a $T_{n}^{\leftarrow,h}$ does not exist, most likely we have ended up in a region of maximal superheating where $S_{3}/T$ approaches a constant (see figure 3). In this case, the action is just too big, since $4\log T/H_{R}$ just falls short of activating nucleation before $T_{\rm max}$. We notice, and this was not present in Buen-Abad et al. (2023), that when the temperature drops below $T_{\rm max}$, the condition for nucleation compares $S_{3}/T$ almost constant and $4\ln(T/H)$, which grows as temperature drops. This implies the possible existence of a solution for a new nucleation can happen with a $T_{n}^{\leftarrow,c}$ while cooling. Eventually there will be another transition $T_{n}^{\rightarrow,c}$ to the minimum in $f$ while cooling. Again, there are either zero or two phase transitions. We show this parameter space in figure 5-right.
As a final comment, we notice that in the region of parameter space where a finite spinodal temperature $T_{\rm SP}$ exists, the inverse transition during heating is guaranteed to occur, and the corresponding direct transition on cooling inevitably follows. However, because the direct cooling transition nucleates at temperatures close to $T_{0}$, the potential barrier lies too much near the origin where perturbation theory is not under control; consequently, at sufficiently low temperatures the exact shape of the potential near $\phi=0$ is uncertain, and one cannot definitively determine the order of this second transition.
@@ V Superheating: inverse transition
The phase transition at high temperature brings the system to a vacuum with a largeN-N number of thermal states. Since they are massless in the new vacuum, they gain momentum when the bubbles expand. Therefore, the leading order contribution from the pressure onto the expanding bubble wall has the opposite sign compared to standard transitions, where the nucleation is towards a minimum with lesser relativistic states Buen-Abad et al. (2023). It has recently been realized Barni et al. (2024, 2025) that this aspect is connected with more generic properties of the hydrodynamics of the plasma surrounding the expanding bubbles, which is treated as a perfect fluid affecting the dynamics of the transition and production of gravitational waves Espinosa et al. (2010); Caprini and others (2019); Hindmarsh et al. (2015, 2017); Giese et al. (2020, 2021); Wang et al. (2022); Ajmi and Hindmarsh (2022); Tenkanen and van de Vis (2022); Wang and Yuwen (2022); Wang et al. (2023).
@@
Here, instead the dynamics is as in Barni et al. (2024, 2025), where the bubbles still nucleate and expand much as in the standard case, but the plasma is flowing inward. In these transitions, the latent heat is negative, and the whole process is endothermic. These transitions are called inverse.
@@
The hallmark of an inverse first‐order transition is that the latent heat is negative, which in the frame of the centre of the bubble corresponds to an inward fluid velocity into the newly nucleated region. A simple diagnostic of this behaviour is the generalised pseudo‐trace of the energy–momentum tensor Barni et al. (2025),
$\alpha_{\vartheta}\equiv{D\vartheta\over 3w_{+}}=\frac{1}{3\,w_{+}}\left[De-\frac{\delta e}{\delta p}Dp\right],$ (19)
@@
and this condition remains valid regardless of the specific equation of state or of the level of approximation. The quantity $\alpha_{\vartheta}$ is nothing but the thermodynamic latent heat, $L=D\vartheta$, computed away from the critical temperature, normalised with the enthalpy of the old phase.
We avoid a full hydrodynamic treatment in favour of the pseudo‐trace criterion, valid when temperature jumps across the wall are small. Recalling that $\bar{\theta}=e-p/c_{s,-}^{2}$ and $\alpha_{\bar{\theta}}=D\bar{\theta}/3w_{+}$ (see Giese et al. (2020)), one finds that the sign of $\alpha_{\bar{\theta}}$ uniquely identifies the transition type.
Figure 6 displays the allowed region in nucleation temperature. When $T_{n}^{\leftarrow}>T_{c}$, these are heating PTs, the latent heat is negative ($\alpha_{\bar{\theta}}^{\leftarrow}<0$), corresponding to inverse transitions—whether the barrier eventually vanishes at a finite spinodal ($T_{c}/T_{0}<3$) or persists indefinitely ($T_{c}/T_{0}>3$). Conversely, for $T_{n}^{\rightarrow}<T_{c}$ the cooling transitions feature $\alpha_{\bar{\theta}}^{\rightarrow}>0$ and exhibit the usual direct behaviour. In these last estimates, we considered $a_{0,2,3}\gg a_{4}$ so that $c_{s}^{2}\simeq 1/3$, appropriate for theories with many relativistic degrees of freedom (cf. Sec. III). The solid red (blue) line indicating the spinodal ($T=T_{0}$) temperature can be regarded as the maximal superheating (supercooling) we can achieve.